# Low-Power Wake-Up Signal Design in 3GPP 5G-Advanced Release 19

### v1

@@ Abstract
The Low-Power Wake-Up Signal (LP-WUS) and Low-Power Synchronization Signal (LP-SS), introduced in 3GPP 5G-Advanced Release 19, represent a major step forward in enabling power-efficient IoT communications. This paper presents a comprehensive overview of the LP-WUS and LP-SS procedures in the RRC_IDLE and RRC_INACTIVE states, and outlines key physical layer design choices. The LP-WUS is designed to be detected by a low-power energy detector (ED), allowing the main radio (MR) to remain switched off. This architecture enables power savings of up to 80% compared to conventional 5G paging mechanisms.
Index Terms:
low-power, 3GPP 5G, wake-up signal
@@ I Introduction
In cellular networks, 3GPP has conducted a Study Item (SI) on LP-WUS in Release 18, cf. [2], with the goal to evaluate the potential reduction in power consumption of a 5G device equipped with a LP Wake-Up Radio (LR). Typically, a 5G device consumes tens of milliwatts even if it is not transmitting or receiving any data, a state called RRC_IDLE/INACTIVE. This idle power consumption is due to the fact that the 5G device has to carry out periodic measurements (to support mobility) and check for potential paging messages (to be reachable), once per Discontinuous Reception (DRX) cycle. Hence, 5G IoT devices with long battery life on a single charge have long been elusive.
@@
The LP-WUS feature takes a step forward and enables significant improvements in power consumption in the idle state while maintaining latency requirements. To achieve this, two signals are introduced, the LP-WUS triggering the decoding of the paging message and LP-SS to perform measurements. Both signals can be processed by the LR which is independent of the 5G Main Radio (MR), i.e. the 5G MR can be powered off while the LP-WUR is active and searching for a potential LP-WUS and performing measurements. This differentiates LP-WUS from existing power-saving features such as the 5G Rel-17 Paging-Early Indication (PEI) which is based on the Physical Data Control CHannel (PDCCH) and requires the MR for reception.
LP-WUS is also specified for RRC_CONNECTED. Similar to RRC_IDLE/INACTIVE, the LP-WUS triggers PDCCH monitoring by the MR. However, since DL traffic arrival is significantly more likely in RRC_CONNECTED, the MR cannot power down into an ultra-deep sleep state since it would take too long to power the MR up again. That is the main reason the power-saving gains are more pronounced in RRC_IDLE/INACTIVE since the MR consumes almost no power in the ultra-deep sleep state. Consequently, this paper focuses on the LP-WUS operation in RRC_IDLE/INACTIVE mode.
The paper is organized as follows: Section II describes the LP-WUS procedures and configurations in RRC_IDLE/INACTIVE mode. In Section III the physical layer design of both LP-WUS and LP-SS is explained. Finally, Section IV provides a summary and conclusion of this contribution.
@@ II-B LP-WUS Occasions and Paging Association
An example of possible PO-to-LO associations is shown in Figure 1. Depending on the configuration, a LO can be associated with 1, 2 or 4 PFs. The configuration $N_{S}>N_{PO}^{LO}$ is (Case D) is supported by indicating multiple values of $T_{PO}^{LO}$ in SIB1, e.g. if $N_{S}=4$ and $N_{PO}^{LO}=1$ there are 4 time offsets signaled because each of the four POs in the PF has its own LO. The network configures $T_{PO}^{LO}$ depending on the wake-up delay of the MR reported by the UEs and the SSB periodicity. For instance, if the UEs in the PO report a wake-up delay of $70ms$, the network configures $T_{PO}^{LO}\geq 70ms$ s.t. there are at least three SSBs available between the LO and PO for measurements and synchronization to receiver the paging message.
@@ II-C Subgroup Signaling in LO
The LP-WUS contains a maximum of $B_{max}=5$ information bits and therefore can, at most, signal 32 values, called codepoints. It is important to retain the possibility to wake up all groups and hence the maximum number of subgroups is 31. Since one LO can be associated to $N_{PO}^{LO}$ POs, the maximum number of subgroups per PO $N_{SG,max}^{PO}\in\{7,15,31\}$ for $N_{PO}^{LO}\in\{1,2,4\}$, respectively. The number of configured subgroups $N_{SG}^{PO}$ is signaled in SIB1.
@@ II-D LP-WUS Occasions within an LO
Each LO can contain $N_{LO}^{MO}\in\{1,2,3,4\}$ LP-WUS Monitoring Occasions (MO) per beam. As described above, it is important to understand that a PO is associated to an LO and not to a MO. By configuring multiple MOs, the network can wake up multiple SGs associated with the LO, because the UE has to monitor all MOs within its LO. For example, consider case C in Figure 1, where with $N_{LO}^{MO}=4$, the network can address SGs in each PO or 4 SGs in a single PO, significantly enhancing paging flexibility.
@@
The duration of a MO, in number of OFDM symbols, is indicated by two values: (i) a nominal MO duration $L_{MO}$ and (ii) an actual MO duration $L$. The difference between the two stems from the number OFDM symbols that are available for LP-WUS transmission. $L$ is the configured duration of the LP-WUS and $L_{MO}$ the duration of the MO in which the LP-WUS is transmitted. If there are not enough OFDM symbols available in the MO, i.e. $L_{MO}<L$, the UE skips this MO, [6].
@@ III Physical Layer Design
One design criteria for LP-WUS is its seamless integration in the existing 5G gNB transmission chain. The Rel-19 LP-WUS and LP-SS support $M\in\{1,2,4\}$ OOK symbols per OFDM symbol which are generated in time-domain and subsequently transformed via DFT into the frequency domain to be mapped to the existing 5G OFDM resource grid.
@@ III-A LP-WUS Design
Let $\mathbf{b}=[b_{0},b_{1},...,b_{B-1}]$ designate the information bits (codepoint) of length $B$ carried by the WUS, where $B\leq 5$ bits. The corresponding time-domain signal $\mathbf{s}=[\mathbf{s}_{0},\mathbf{s}_{1},...,\mathbf{s}_{L-1}]$ spans $L$ OFDM symbols (not necessarily consecutive). Each of the OOK waveforms within an OFDM symbol $\mathbf{s}_{l}=[\mathbf{r}_{0},\mathbf{r}_{1},...,\mathbf{r}_{M-1}]$ is constituted of $M$ OOK symbols. Thus, $\mathbf{s}$, $\mathbf{s}_{l}$ and $\mathbf{S}_{l}$ denote the time-domain OOK signals of the entire WUS transmission, the WUS per OFDM symbol and the frequency-domain representation, respectively.
@@ III-A1 Encoding
The encoding procedure consists of three steps: (i) Channel coding, (ii) rate-matching (RM) followed by (iii) line coding. Channel coding is necessary to meet the requirements on False-Alarm Rate (FAR) and Missed-Detection Rate (MDR) and uses the coding schemes for small block lengths defined in [4, Section 5.3.3] with $Q_{m}=1$. The $N$ coded bits $d_{i}$, $i=0,1,...,N-1$ with $N=1$, $N=3$ and $N=32$ for $B=1$, $B=2$ and $B>2$, respectively, are given by
@@
where the base sequences $M_{i,k}$ of the Reed-Muller code are defined in [4, Table 5.3.3.3-1].
Subsequently, rate matching is applied to the coded bits $d_{i}$ to obtain a rate-matched bit sequence $f_{k}$, $k=0,1,...,E-1$, [4, Section 5.4.3]
@@
In the last step, the well-known Manchester Code (MC) is applied to $f_{k}$ resulting in $g_{m}$, $m=0,1,...,G-1$ coded bits. The rate $R=1/2$ MC maps a single input bit to two coded bits, i.e.
@@
with $k=0,1,...,E-1$. The main advantage of Manchester coding is that it allows for a simple and robust decoder. More precisely, the decoder simply compares a metric (e.g. received energy) corresponding to the two encoded bits as opposed to a threshold which is difficult to obtain in fading channels.
@@ III-A2 LP-WUS Modulation
where $\mathbf{r}_{c_{m}}$ is the ON-sequence $c_{m}$, $c_{m}\in\{0,1,...,N_{seq}-1\}$ denotes the sequence index of the $m^{th}$ OOK ON-symbol ($m=0,1,...,E-1$) and $N_{seq}$ is the number of configured sequences. The WUS time-domain signal $\mathbf{s}_{l}$ for OFDM symbol $l$ is the concatenation of the $M$ ON-sequences $\mathbf{r}$.
@@ III-A3 Wake-up Sequence Design
The design of the ON-sequence $\mathbf{r}$, mapped to one OOK symbol of length ${M_{\rm ZC}}$, is subject to the following criteria: (i) Good spectral properties (e.g. low PAPR), (ii) low specification effort (i.e. reuse of existing sequences) and (iii) good correlation properties.
The sequences in [5, TS 38.211] are reused and known as cyclically extended Zadoff-Chu sequences, [10]. The $c^{th}$ ON-sequence $\mathbf{r}_{c}$ reads
@@
where $c=0,1,...,N_{seq}-1$. For a fixed LP-WUS bandwidth of $N_{SC}^{WUS}=132$ subcarriers, the length ${M_{\rm ZC}}=N_{SC}^{WUS}/M\in\{132,66,33\}$ for $M\in\{1,2,4\}$, respectively. How the ON-sequences are chosen is described in the next section.
@@ III-A4 Sequence Encoding
Multiple sequences can be configured to allow a coherent receiver to obtain the information bits $\mathbf{b}$ through sequence correlation. The maximum number of sequences allowed $N_{seq}^{max}$ depends on $M$ as $N_{seq}^{max}=\{16,8,4\}$ for $M=\{1,2,4\}$, respectively. The reason is that a longer sequence (smaller values of $M$) supports more sequences with good correlation properties than a short sequence.
@@
with $i=0,1,...,E_{s}-1$, where $E_{s}=E\delta$ and $E=LM/2$ is the number of OOK ON-symbols available. Segment $\mathbf{f}_{s}$ into blocks of $\delta$ bits such that $\mathbf{f}_{s}=[\mathbf{f}_{s,0},\mathbf{f}_{s,1},...,\mathbf{f}_{s,(E-1)}]$ where block $m=0,1,...,E-1$ is given by
@@ 3 $[101001]$ $[011001011001]$ $[1010100110100110]$
TABLE II: Binary LP-SS $\mathbf{b}_{LPSS}$ of the short sequences.
@@ III-C Receiver Design
with $k=0,1,...,G-1$. The estimate $\hat{\mathbf{f}}$ is then passed to the channel decoder to obtain the payload estimate $\hat{\mathbf{b}}$. A different implementation can also directly correlate the received envelope with all possible encoded messages resulting in a higher processing gain and improved performance.
@@
IV Conclusion
This tutorial paper provides a comprehensive overview of the LP-WUS feature for the RRC_IDLE/INACTIVE states in Release 19, including a detailed discussion of the physical layer design of both LP-WUS and the LP-SS. Energy efficiency is a key objective in 6G, and a single LP-WUS-based UE power saving mechanism is a strong candidate for future 6G systems, as it can be received by both low-power energy detectors and conventional coherent receivers.

### v2

@@ Abstract
The Low-Power Wake-Up Signal (LP-WUS) and Low-Power Synchronization Signal (LP-SS), introduced in 3GPP 5G-Advanced Release 19, mark an important advancement toward power-efficient IoT communications. This paper provides a comprehensive overview of the LP-WUS procedures in the RRC_IDLE and RRC_INACTIVE states and summarizes the key physical-layer design aspects. The LP-WUS is intended to be detected by a low-power energy detector (ED), allowing the main radio (MR) to remain switched off, thereby enabling substantial power savings compared to conventional 5G paging mechanisms. As such, LP-WUS is considered the baseline for 6G WUS design. Furthermore, different receiver architectures are evaluated, highlighting the inherent trade-offs between power-saving gains and coverage performance.
@@ I Introduction
In cellular networks, 3GPP has conducted a Study Item (SI) on LP-WUS in Release 18, cf. [2], [4], with the goal to evaluate the potential reduction in power consumption of a 5G device equipped with a LP Wake-Up Radio (LR). Typically, a 5G device consumes tens of milliwatts even if it is not transmitting or receiving any data, a state called RRC_IDLE/INACTIVE. This idle power consumption is due to the fact that the 5G device has to carry out periodic measurements (to support mobility) and check for potential paging messages (to be reachable), once per Discontinuous Reception (DRX) cycle. Hence, 5G IoT devices with long battery life on a single charge have long been elusive.
@@
The LP-WUS feature takes a step forward and enables significant improvements in power consumption in the idle state while maintaining latency requirements. To achieve this, two novel signals are introduced, the LP-WUS triggering the decoding of the paging message and the LP-SS to perform measurements. Both signals are processed by the LR, operating independently of the 5G Main Radio (MR). In this way, the MR can remain powered down while the LP-WUR continues to search for potential LP-WUS and carry out measurements via LP-SS, thereby reducing energy consumption. This differentiates LP-WUS from existing power-saving features such as the 5G Rel-17 Paging-Early Indication (PEI) which is based on the Physical Data Control CHannel (PDCCH) and requires the MR for reception.
LP-WUS is also specified for the RRC_CONNECTED mode. Similar to RRC_IDLE/INACTIVE, the LP-WUS triggers PDCCH monitoring by the MR. However, since DL traffic arrival is significantly more likely in RRC_CONNECTED, the MR cannot power down into an ultra-deep sleep state since it would take too long to power the MR up again. That is the main reason the power-saving gains are more pronounced in RRC_IDLE/INACTIVE since the MR consumes almost no power in the ultra-deep sleep state. Consequently, this paper focuses on the LP-WUS operation in RRC_IDLE/INACTIVE mode.
The paper is organized as follows: Section II describes the LP-WUS procedures and configurations in RRC_IDLE/INACTIVE mode. In Section III the physical layer design of both LP-WUS and LP-SS is explained. Finally, Section V provides a summary and conclusion of this contribution.
@@ II-B LP-WUS Occasions and Paging Association
An example of possible PO-to-LO associations is shown in Figure 1. Depending on the configuration, a LO can be associated with 1, 2 or 4 PFs. The configuration $N_{S}>N_{PO}^{LO}$ (Case D) is supported by indicating multiple values of $T_{PO}^{LO}$ in SIB1, e.g. if $N_{S}=4$ and $N_{PO}^{LO}=1$ there are 4 time offsets signaled because each of the four POs in the PF has its own LO. The network configures $T_{PO}^{LO}$ depending on the wake-up delay of the MR reported by the UEs and the SSB periodicity. For instance, if the UEs with the same PO report a wake-up delay of $70ms$, the network configures $T_{PO}^{LO}\geq 70ms$ s.t. there are at least three SSBs available between the LO and PO for measurements and synchronization to receive the paging message with the MR.
@@ II-C Subgroup Signaling in LO
The LP-WUS contains a maximum of $B_{max}=5$ information bits and therefore can, at most, signal 32 values, called codepoints. It is important to retain the possibility to wake up all groups and hence the maximum number of subgroups is 31. Since one LO can be associated to $N_{PO}^{LO}$ POs, the maximum number of subgroups per PO is $N_{SG,max}^{PO}\in\{31,15,7\}$ for $N_{PO}^{LO}\in\{1,2,4\}$, respectively. The number of configured subgroups $N_{SG}^{PO}$ is signaled in SIB1.
@@ II-D LP-WUS Occasions within an LO
Each LO can contain $N_{LO}^{MO}\in\{1,2,3,4\}$ LP-WUS Monitoring Occasions (MO) per beam. As described above, it is important to understand that a PO is associated to an LO and not to a MO. By configuring multiple MOs, the network can wake up multiple SGs associated with the LO, because the UE has to monitor all MOs within its LO. For example, consider case C in Figure 1, assuming $N_{LO}^{MO}=4$, the network can either address SGs in each PO or 4 SGs in a single PO, significantly enhancing paging flexibility.
@@
The duration of a MO, in number of OFDM symbols, is indicated by two values: (i) a nominal MO duration $L_{MO}$ and (ii) an actual MO duration $L$. The difference between the two stems from the number OFDM symbols that are available for LP-WUS transmission. $L$ is the configured duration of the LP-WUS and $L_{MO}$ the duration of the MO in which the LP-WUS is transmitted. If there are not enough OFDM symbols available in the MO, i.e. $L>L_{MO}$, the UE skips this MO, [7].
@@ III Physical Layer Design
One design criteria for LP-WUS is its seamless integration in the existing 5G gNB transmission chain. The Rel-19 LP-WUS and LP-SS support $M\in\{1,2,4\}$ OOK symbols per OFDM symbol which are generated in time-domain and subsequently transformed via DFT into the frequency domain to be mapped to the existing 5G OFDM resource grid. For a fixed LP-WUS signal bandwidth (excluding guard-bands) of $N_{SC}^{WUS}=132$ subcarriers, the length of the OOK sequences are ${M_{\rm ZC}}=N_{SC}^{WUS}/M\in\{132,66,33\}$ for $M\in\{1,2,4\}$, respectively.
@@ III-A LP-WUS Design
Let $\mathbf{b}=[b_{0},b_{1},...,b_{B-1}]$ designate the information bits (codepoint) of length $B$ carried by the WUS, where $B\leq 5$ bits. The corresponding time-domain signal $\mathbf{s}=[\mathbf{s}_{0},\mathbf{s}_{1},...,\mathbf{s}_{L-1}]$ of length $N_{SC}^{WUS}L$ spans $L$ OFDM symbols (not necessarily consecutive). Each of the OOK waveforms within an OFDM symbol $\mathbf{s}_{l}=[\mathbf{r}_{0},\mathbf{r}_{1},...,\mathbf{r}_{M-1}]$ is constituted of $M$ OOK symbols. Thus, $\mathbf{s}$, $\mathbf{s}_{l}$ and $\mathbf{S}_{l}$ denote the time-domain OOK signals of the entire WUS transmission, the WUS per OFDM symbol and the frequency-domain representation, respectively.
@@ III-A1 Encoding
The encoding procedure consists of three steps: (i) Channel coding, (ii) rate-matching (RM) followed by (iii) line coding. Channel coding is necessary to meet the requirements on False-Alarm Rate (FAR) and Missed-Detection Rate (MDR) and uses the coding schemes for small block lengths defined in [5, Section 5.3.3] with $Q_{m}=1$. The $N$ coded bits $d_{i}$, i=0,1,,N1i=0,1,...,N-1 with $N=1$, $N=3$ and $N=32$ for $B=1$, $B=2$ and $B>2$, respectively, are given by
@@
where the base sequences $M_{i,k}$ of the Reed-Muller code are defined in [5, Table 5.3.3.3-1].
Subsequently, rate matching is applied to the coded bits $d_{i}$ to obtain a rate-matched bit sequence $f_{k}$, k=0,1,,E1k=0,1,...,E-1, [5, Section 5.4.3]
@@
In the last step, the well-known Manchester Code (MC) is applied to $f_{k}$ resulting in $g_{m}$, m=0,1,,G1m=0,1,...,G-1 coded bits. The rate $R=1/2$ MC maps a single input bit to two coded bits, i.e.
@@
with k=0,1,,E1k=0,1,...,E-1. The main advantage of Manchester coding is that it allows for a simple and robust decoder. More precisely, the decoder simply compares a metric (e.g. received energy) corresponding to the two encoded bits as opposed to a threshold which is difficult to obtain in fading channels.
@@ III-A2 LP-WUS Modulation
where $\mathbf{r}_{c_{m}}$ is the ON-sequence $c_{m}$, $c_{m}\in\{0,1,...,N_{seq}-1\}$ denotes the sequence index of the $m^{th}$ OOK ON-symbol (m=0,1,,E1m=0,1,...,E-1) and $N_{seq}$ is the number of configured sequences. The WUS time-domain signal $\mathbf{s}_{l}$ for OFDM symbol $l$ is the concatenation of the $M$ ON-sequences $\mathbf{r}$.
@@ III-A3 Wake-up Sequence Design
The design of the ON-sequence $\mathbf{r}_{c_{m}}$, mapped to one OOK symbol of length $M_{\rm ZC}$}, is subject to the following criteria: (i) Good spectral properties (e.g. low PAPR), (ii) low specification effort (i.e. reuse of existing sequences) and (iii) good correlation properties.
The sequences in [6, TS 38.211] are reused and known as cyclically extended Zadoff-Chu sequences, [11]. The $c^{th}$ ON-sequence $\mathbf{r}_{c}$ reads
@@
where $c=0,1,...,N_{seq}-1$. How the ON-sequences are chosen is described in the next section.
@@ III-A4 Sequence Encoding
Multiple sequences can be configured to allow a coherent receiver (i.e. with phase information) to obtain the information bits $\mathbf{b}$ through sequence correlation. The maximum number of sequences allowed $N_{seq}^{max}$ depends on $M$ as $N_{seq}^{max}=\{16,8,4\}$ for $M=\{1,2,4\}$, respectively. The reason is that a longer sequence (smaller values of $M$) supports more sequences with good correlation properties than a short sequence.
@@
with $i=0,1,...,E_{s}-1$, where $E_{s}=E\delta$ and $E=LM/2$ is the number of OOK ON-symbols available. Segment $\mathbf{f}_{s}$ into blocks of $\delta$ bits such that $\mathbf{f}_{s}=[\mathbf{f}_{s,0},\mathbf{f}_{s,1},...,\mathbf{f}_{s,(E-1)}]$ where block m=0,1,,E1m=0,1,...,E-1 is given by
@@ 3 $[101001]$ $[011001011001]$ $[1010100110100110]$
TABLE II: Binary LP-$S$ $\mathbf{b}_{LPSS}$ of the short sequences.
@@ III-C Receiver Design
with k=0,1,,G1k=0,1,...,G-1. Alternatively, to improve decoding performance, softbits for $\hat{f}_{k}$ can be obtained by computing the difference $e_{2k}-e_{2k+1}$. The estimate $\hat{\mathbf{f}}$ is then passed to the channel decoder to obtain the payload estimate $\hat{\mathbf{b}}$. A different implementation can also directly correlate the received envelope with all possible encoded messages resulting in a higher processing gain and improved performance.
@@
III-D Comparison to IEEE 802.11ba
The Wi-Fi standard IEEE 802.11ba [1] also specifies an LP-WUS. Since both 5G and Wi-Fi are based on CP-OFDM, comparing the respective design choices is of particular interest. A first notable difference is that Wi-Fi does not support mobility; therefore, a dedicated signal for low-power measurements such as the LP-SS is not required. In 3GPP, the LP-SS also serves as a synchronization reference for LP-WUS reception, whereas Wi-Fi uses a preamble preceding the WUS for the same purpose. Although a preamble-based solution was considered in 3GPP, it was eventually deemed unnecessary, as synchronization performance can instead be controlled by adjusting the LP-SS periodicity.
Another key difference lies in the subcarrier spacing. Wi-Fi uses an SCS of 312.5 kHz, which is significantly higher than in 3GPP; this results in very short OFDM symbol durations. Consequently, Wi-Fi does not need to multiplex multiple OOK symbols within one OFDM symbol, and therefore uses $M=1$. However, two data rates are supported in Wi-Fi: the higher rate is achieved by using only every second subcarrier, effectively halving the symbol duration. Both systems employ Manchester coding, but only 3GPP applies channel coding to enhance coverage. Additionally, Wi-Fi does not specify the ON-sequence (BPSK and QAM modulation are used), and therefore the LP-WUS in Wi-Fi can be received only with an envelope detector.
III-E WUS Design in 6G
Energy efficiency is one of the key components of 6G and on the UE side, a unified WUS design is under discussion. One major issue with the LP-WUS in 5G is coverage, because a paging message cannot be missed due to non-detection of the WUS. An ED detector has a significantly reduced coverage than a CD and also requires the LP-SS for synchronization and measurements. A WUS design based on CD eliminates these drawbacks since the CD has better performance and can use the legacy SSB for synchronization and measurements. Hence, the current discussions move away from OOK modulation and favor designs based on coherent detection.
IV Numerical Evaluation
In this section, we evaluate the performance in terms of BLER, i.e. if the entire WUS payload is received correctly or not, for three different receiver types. The ED accumulating energy per OOK symbol and computing softbits for channel decoding, A CD-S, performing CD through sequence correlation to obtain $B$ in the shortest amount of time and CD-L which correlates over the entire WUS transmission.
Parameter Value
Carrier Frequency 2.6 GHz
Sub-carrier Spacing 30 kHz
System BW 20 MHz (51 PRBs)
Antenna Config SISO
WUS BW 5 MHz (132 SCs)
WUS Sampling Rate 7.68 MHz
LP Filter 3rd order Butterworth
LP Filter BW 4.32 MHz
Channel Model TDL-C 300ns
TABLE III: Simulation assumption.
Figure 7 shows the BLER performance for $B=5$ for the different receivers. Overall, it can be seen that CD performs significantly better than ED. As expected, CD-L has the best performance independent of the number of sequences used. On the other hand, the performance of CD-S depends on $N_{seq}$, i.e. for $N_{seq}=2$, CD-S requires 5 OFDM symbols to recover the payload $B$ whereas for $N_{seq}=8$, only 2 OFDM symbols are required. This results in a 5dB SNR loss at 1%BLER, because less energy is captured and the cross-correlation among the sequences increased. This highlights the trade-off for CD, shorter detection saves power at the expense of detection performance. A practical receiver, will likely be hybrid, i.e. employing ED in good channel conditions and switching to CD if conditions get worse.
[…]
Figure 7: Performance comparison, $B=5$, $M=2$, $L=14$, Table III.
V Conclusion
This tutorial paper provides a comprehensive overview of the LP-WUS feature for the RRC_IDLE/INACTIVE states in Release 19, including a detailed discussion of the physical layer design of both LP-WUS and the LP-SS. Energy efficiency is a key objective for 6G, and developing a unified WUS design that achieves an optimal balance between power-saving gains and coverage remains a key challenge to be addressed in 6G.